The perimeter of a rectangular swimming pool is 70 meters. The pool is twice as long as it is wide. What is the length of the pool?
Answer by lwsshak3(11628)

The perimeter of a rectangular swimming pool is 70 meters. The pool is twice as long as it is wide. What is the length of the pool?
***
let x=width of pool
2x= length of pool
2(width+length)=perimeter
2(x+2x)=70
6x=70
x=70/6
2x=70/3=23-1/3
What is the length of the pool? 23-1/3 m